OpenStudy (anonymous):

4x-14=10 how do i solve this equation? show me in steps plz.

OpenStudy (pokemon23):

add 14 to 10

OpenStudy (saifoo.khan):

4x-14=10 add 14 on both sides, 4x = 10 + 14 4x = 24 divide 4 from both sides, x = 24/4 x = 6

OpenStudy (pokemon23):

grrr nj

OpenStudy (anonymous):

What Saifoo said  Oh yeah, who likes Charlie Brown?